# Roadmap for electronic structure, anharmonicity, and electron-phonon calculations in locally disordered inorganic and hybrid halide perovskites

### v1

@@ V.2 LO-TO splitting
The LO-TO splitting associated with the high energy LO mode is the smallest in FA-based perovskites. For example, the LO-TO splitting in FAPbI3 is 3.6 meV, in MAPbI3 is 5.8 meV, and in CsPbI3 is 4.9 meV. To rationalize this result we first use the simplified expression for the non-analytic LO-TO splitting, $\Delta_{\rm LO-TO}\propto|Z^{*}|^{2}/L^{3}\epsilon^{\infty}$, where $Z^{*}$ indicates the Born effective charges in units of elementary charge, $\epsilon^{\infty}$ is the high frequency dielectric constant, and $L^{3}$ is the volume of the cubic system. In view of the expression for $\Delta_{\rm LO-TO}$, one can expect FAPbI3 to exhibit a similar if not larger splitting among the A-site variants. This is based on the fact that FAPbI3 has relatively larger Born effective charges (e.g., $Z_{\rm Pb}^{*}=4.72$, $Z_{\rm I}^{*}=4.68$), a similar high-frequency dielectric constant ($\epsilon^{\infty}=6.13$), and a similar lattice constant ($L=6.36$) compared to MAPbI3 (with $Z_{\rm Pb}^{*}=4.59$, $Z_{\rm I}^{*}=4.44$, $\epsilon^{\infty}=5.87$, and $L=6.31$) and CsPbI3 (with $Z_{\rm Pb}^{*}=4.49$, $Z_{\rm I}^{*}=4.17$, $\epsilon^{\infty}=5.75$, and $L=6.25$). Nonetheless, our density functional perturbation theory calculations show that FAPbI3 exhibits the smallest LO-TO splitting among the three compounds. This contradiction indicates that the simplified expression does not capture the full microscopic picture Gonze and Lee (1997), as it neglects the phonon mode character, dipole directionality, and mode mixing. We ascribe the reduced LO-TO splitting in FAPbI3 to enhanced mode mixing between polar optical phonons and internal molecular vibrations [Fig. 13(c)], as well as directional coupling effects that diminish the effective dipole strength of the LO mode.
V.3 Overdamped phonon dynamics
@@
V.4 Thermal induced renormalization
We have seen that thermal induced corrections are strongly modified for most cubic compounds in the presence of local disorder. As shown in Fig. 10(a), electron-phonon contribution is strongly suppressed for all compounds, with the largest reductions found for Cs-based compounds which exhibit the largest degree of local disorder. The only case where electron-phonon coupling contributes negatively to the temperature coefficient of the band gap leading to a band gap closing is MAPbCl3. We attribute this to the fact that MAPbCl3 exhibits one of the highest degrees of local disorder among the hybrid halide perovskites, reaching levels comparable to those of Cs-based compounds [Fig. 2(b) and Fig. 2(c) of Ref. [Zacharias et al., 2025]]. The opposing effects of electron-phonon coupling and thermal expansion in polymorphous cubic MAPbCl3 lead to the nearly constant variation of the band gap with temperature, consistent with experiments.
@@ V.4 Thermal induced renormalization
Our results for the total temperature coefficients of the band gap using ZG polymorphous structures compare well with the experimental values, as shown quantitatively in Fig. 10(c) and Figs. 9(a)-(d), (i)-(l), and (q)-(t). Importantly, we have seen that even using a single ZG polymorphous configuration obtained from the MSS structure is a good approximation for capturing the temperature variation of the band gap. We also find that thermal-induced corrections lead to a linear variation of the effective masses with temperature [Figs. 9(e)-(h), (m)-(p), and (u)-(x)]. In APbI3 and APbBr3 compounds, the hole effective masses remain larger than the electron effective masses, whereas in ASnI3 and APbCl3 compounds, electron effective masses remain larger than hole effective masses. We note that our finite temperature calculations: (i) rely on a constant temperature phonon approximation, (ii) neglect non-adiabatic effects in electron-phonon renormalization of the band structure Poncé et al. (2015); Miglio et al. (2020); Engel et al. (2022), (iii) neglect many-body quasiparticle corrections to the electronic structure Antonius et al. (2014); Marini et al. (2015); Li et al. (2024), and (iv) omit phonon renormalization due to thermal expansion. While including these effects may improve quantitative agreement with experiment, we do not expect them to alter the qualitative trends, as our current results already show good overall agreement.
@@ VI.1 Electronic structure calculations
DFT calculations were performed using Quantum Espresso (QE) Giannozzi et al. (2009, 2017). We employed optimized norm-conserving Vanderbilt pseudopotentials from the PseudoDojo library Hamann (2013); van Setten et al. (2018) and the Perdew-Burke-Ernzerhof exchange-correlation functional revised for solids (PBEsol) Perdew et al. (2008). The lattice constants ($L$) of all cubic phases, reported in Table 1, were fixed close to their experimental values Mashiyama et al. (2003); Rakita et al. (2015); Whitfield et al. (2016); Dang et al. (2016); Schueller et al. (2017); Kontos et al. (2018); Marronnier et al. (2018); Handa et al. (2020); He et al. (2021). Calculations for the monomorphous structures (5 atoms for inorganic and 12 atoms for hybrid halide perovskites) were performed using a plane wave cutoff energy of 120 Ry and $6\times 6\times 6$ uniform ${\bf k}$-grids. For exploring reference and polymorphous structures with the ASDM as described in Sec. II.3, we performed geometry optimization in $2\times 2\times 2$ supercells (40 atoms for inorganic and 96 atoms for hybrid halide perovskites), until residual forces per atom were below 3$\times$10-4 eV/Å  using a plane wave cutoff energy of 120 Ry and $3\times 3\times 3$ uniform ${\bf k}$-grids. During geometry optimization the lattice constants were kept fixed. We note that for the reference structures of hybrid halide perovskites only the molecules were relaxed, while for the unique reference (monomorphous) structures of inorganic compounds no relaxation was performed as the residual forces were zero. For each inorganic halide perovskite, we generated 1 high-symmetry monomorphous structure and 10 polymorphous structures. For each hybrid halide perovskite, we generated 1 MSS reference structure and its corresponding polymorphous structure, as well as 10 reference structures obtained by relaxing molecules from random initial orientations and their corresponding polymorphous structures. To construct MSS structures for hybrid halide perovskites, we employed a $2\times 2\times 2$ supercell and initially fixed the orientation of one molecular cation within a single unit cell. Under this supercell constraint, the remaining seven molecules were placed such that a subset of the original space group symmetries, corresponding to the P422 space group, was preserved. The band gap of each structure was obtained by including SOC effects via fully relativstic pseudopotentials and evaluating the energy difference between the conduction band minimum and valence band maximum at the R point.
@@
PBE0 hybrid functional calculations Perdew et al. (1996); Adamo and Barone (1999) for the reference and polymorphous structures were performed using the VASP code Kresse and Furthmüller (1996) within the projector augmented-wave formalism Kresse and Joubert (1999). A plane-wave energy cut-off of 400 eV was used, and the Brillouin zone was sampled with a 2$\times$2$\times$2 uniform k-grid.
@@ VI.3 Diffuse scattering
One-phonon diffuse scattering maps of MAPbBr3, shown in Fig. 6, were computed using disca.x of EPW Zacharias et al. (2021a, b); Lee et al. (2023). The Debye-Waller factors at each scattering wavevector ${\bf Q}$ were evaluated for a 30$\times$30$\times$30 ${\bf q}$-grid using phonon polarization vectors and frequencies obtained by Fourier interpolation of the effective IFCs computed at 300 K with ASDM for 2$\times$2$\times$2 supercells. A 30$\times$30$\times$1 uniform ${\bf Q}$-grid per Brillouin zone together with the computed Debye-Waller factors were used to calculate the one-phonon diffuse scattering intensity in reciprocal lattice planes perpendicular to the Cartesian $z$-axis. Parameters from Ref. [Peng et al., 2004] were used to model the atomic scattering amplitudes as a sum of Gaussians.
@@ VI.4 Effective masses
Effective masses were computed at the DFT level including SOC and by evaluating the diagonal components of the effective mass tensor via finite differences Giustino (2014), followed by an isotropic average. That is we used the second-order finite difference formula based on the energies at the R point and nearby ${\bf k}$-points ($|\Delta{\bf k}|$ = 0.01 $2\pi$/$\alpha$) along each Cartesian direction. The effective mass enhancement factors due to polymorphism were determined as $\lambda=m^{*}_{\rm poly}/m^{*}_{\rm ref}-1$ and due to thermal effects as $\lambda(T)=[m^{*}+\Delta m^{*}(T)]/m^{*}-1$. The reduced effective masses were evaluated as $\mu=m^{*}_{\rm h}m^{*}_{\rm e}/(m^{*}_{\rm h}+m^{*}_{\rm e})$. The slopes describing the temperature variation of the effective masses are denoted by $\lambda^{T}$.
@@ VI.5 Electron-phonon calculations
To account for electron-phonon coupling effects at various temperatures, we employed special displacements (or ZG displacements) Zacharias and Giustino (2020) generated via Eq. (6) for 4$\times$4$\times$4 supercells. We considered a constant temperature approximation for the phonons, i.e. anharmonic phonons computed at a specific temperature for each material (specified in Fig. 8), and the resulting vibrational modes were thermally populated at various temperatures to generate special displacements. We applied the special displacements on the nuclei of both the reference and polymorphous structures leading to the ZG reference and ZG polymorphous structures. We note that the initial 4$\times$4$\times$4 reference and polymorphous structures (containing 320 atoms for inorganic and 768 atoms for hybrid halide perovskites) were constructed by replicating the corresponding structures obtained in 2$\times$2$\times$2 supercells. SOC was included in all electron-phonon coupling calculations. To reduce computational cost, the plane wave energy cutoff was set to 60 Ry. Convergence tests performed on a subset of materials (MAPbBr3, FAPbBr3, FAPbI3, and MASnI3) using 2$\times$2$\times$2 supercells showed that the band gap values differ by less than 1 meV compared to calculations with a 120 Ry cutoff. Furthermore, the Brillouin zone of the 4$\times$4$\times$4 supercell was sampled using a 1$\times$1$\times$1 ${\bf k}$-grid. In this case, we found for the same subset of materials that the band gap varied by less than 10 meV and the thermal induced renormalization by less than 1 meV.
To determine the contribution of electron-phonon coupling to the band gap renormalization and effective masses ($dE_{\rm g}/dT|_{\rm EPC}$ and $dm^{*}/dT|_{\rm EPC}$), we applied linear fits to the data points of each material computed at various temperatures without considering the effect of thermal expansion. Each data point was obtained as a Boltzmann-weighted average [Eq. (11)] over the band gap of 10 ZG reference or 10 ZG polymorphous structures.
@@ VI.6 Thermal expansion contribution
To evaluate the contribution of thermal expansion to the temperature dependence of the band gap, $dE_{\rm g}/dT|_{\rm TE}$, and effective masses, $dm^{*}/dT|_{\rm TE}$, we employed the following strategy for each material: (i) The lattice constants of the initial 10 reference and 10 polymorphous 2$\times$2$\times$2 supercells were varied by $\Delta L$ = $\pm$0.04 Å. (ii) For each modified lattice, geometry optimizations were performed while keeping the lattice constants fixed, yielding new reference and polymorphous configurations. (iii) Band gaps and effective masses were computed for each individual structure, and averaged over the set of configurations corresponding to the same lattice constant. (iv) The derivatives $dE_{\rm g}/dL$ and $dm^{*}/dL$ were estimated by fitting a linear function to the computed averaged values at $L-\Delta L$, $L$, and $L+\Delta L$, equivalent to a central finite difference approximation. (v) Experimental values of the linear thermal expansion coefficient, $\alpha=1/L\,dL/dT$, of each material were taken from the literature Mashiyama et al. (2003); Rakita et al. (2015); Whitfield et al. (2016); Dang et al. (2016); Schueller et al. (2017); Kontos et al. (2018); Marronnier et al. (2018); Handa et al. (2020); He et al. (2021). Due to the absence of experimental data for FAPbCl3, we used $\alpha=0.6\times 10^{-4}$ K-1 close to the value reported for MAPbCl3 Handa et al. (2020). (vi) Finally, the thermal expansion contributions for each material were evaluated using the relations:
@@ VI.7 Phonon spectral functions
Phonon spectral functions of the polymorphous structures shown in Fig. 11 were obtained by phonon unfolding using the approach described in Ref. [Zacharias et al., 2023a] and the ZG.x code of EPW Lee et al. (2023). We used the IFCs calculated for the polymorphous structures without enforcing any translational or rotational symmetries of the unit cell or the reference structures. Based on the computed eigenvectors, which reflect spatial correlations introduced by local disorder, we determined the spectral weights of each configuration using 394 equally-spaced ${\bf q}$-points along the the X-R-M-$\Gamma$-R path of the Brillouin zone. Convergence of the spectral weights was achieved by using a 12$\times$12$\times$12 ${\bf g}$-grid of reciprocal lattice vectors.
@@ VII Conclusions and outlook
Based on the ASDM, we developed a unified methodology for electronic structure, anharmonicity, and electron-phonon coupling calculations in locally disordered inorganic and hybrid halide perovskites. We carried out a comparative high-throughput study of 12 cubic compounds and analyzed the modifications due to local disorder to their electronic structures with and without thermal effects. In addition to generating polymorphous structures, we also investigated reference configurations obtained by relaxing only the molecular sublattice, in order to minimize artifacts arising from net dipole moments, particularly relevant for MA-based compounds, and to enable systematic comparisons with the fully polymorphous structures. Our calculations reveal strong correlations between band gap widening and effective mass enhancement with structural distortions, such as changes in bond angles and lengths, induced by positional polymorphism. We further demonstrate that the extent of positional polymorphism is strongly influenced by the size of the A-site cation, a trend that correlates with the tolerance factor Zacharias et al. (2025), where lower values typically indicate increased structural flexibility and a greater propensity for local distortions.
@@
We evaluated the temperature dependence of band gaps and effective masses using both ZG reference structures and ZG polymorphous structures for all cubic compounds. For the band gaps, we achieved excellent agreement with experimental values, highlighting the importance of incorporating corrections due to positional polymorphism, exchange-correlation treatments beyond standard DFT, anharmonic electron-phonon coupling, and thermal expansion in order to accurately reproduce experimental trends. We further quantified the contributions of electron-phonon coupling and thermal expansion to the temperature-induced band gap renormalization, showing that electron-phonon coupling dominates in Cs-based and MA-based compounds, whereas thermal expansion plays the leading role in FA-based systems. This behavior is attributed to the larger volumetric thermal expansion coefficients of FA-based compounds. In the case of locally disordered FASnI3, we found that the stereochemically active Sn lone pair enhances the thermal expansion contribution, which primarily accounts for the largest temperature coefficient of the band gap among the materials studied. Finally, we showed that even a single ZG polymorphous configuration, generated from the MSS structure and ASDM phonons, yields results comparable to Boltzmann-weighted averages over 10 ZG polymorphous configurations.
@@
To our knowledge, this is the first report of anharmonic quasiparticle phonon dispersions in hybrid halide perovskites, validated against experimental measurements. The anharmonic, temperature-dependent interatomic force constants (IFCs), along with the corresponding reference and polymorphous structures, which will be made available in NOMAD, provide a foundation for state-of-the-art perturbative electron-phonon calculations Baroni et al. (2001); Giustino (2017); Sangalli et al. (2019); Gonze et al. (2020); Engel et al. (2020); Zhou et al. (2021); Yang et al. (2022); Lee et al. (2023); Li et al. (2024) in cubic hybrid halide perovskites. These calculations can help elucidate key phenomena such as charge carrier mobilities Poncé et al. (2019); Ranalli et al. (2024), polaron formation Lafuente-Bartolome et al. (2024), as well as non-equilibrium dynamics Caruso (2021). Furthermore, our finding that a single ZG polymorphous configuration within a $4\times 4\times 4$ supercell provides an excellent approximation for nonperturbative electron-phonon calculations holds promise for extending this approach to large-scale, high-throughput studies across other hybrid halide perovskite systems.

### v2

@@ V.3 LO-TO splitting
The LO-TO splitting associated with the high energy LO mode is the smallest in FA-based perovskites. For example, the LO-TO splitting in FAPbI3 is 3.6 meV, in MAPbI3 is 5.8 meV, and in CsPbI3 is 4.9 meV. To rationalize this result we first use the simplified expression for the non-analytic LO-TO splitting, $\Delta_{\rm LO-TO}\propto|Z^{*}|^{2}/L^{3}\epsilon^{\infty}$, where $Z^{*}$ indicates the Born effective charges in units of elementary charge, $\epsilon^{\infty}$ is the high frequency dielectric constant, and $L^{3}$ is the volume of the cubic system. In view of the expression for $\Delta_{\rm LO-TO}$, one can expect FAPbI3 to exhibit a similar if not larger splitting among the A-site variants. This is based on the fact that FAPbI3 has relatively larger Born effective charges (e.g., $Z_{\rm Pb}^{*}=4.72$, $Z_{\rm I}^{*}=4.68$), a similar high-frequency dielectric constant ($\epsilon^{\infty}=6.13$), and a similar lattice constant ($L=6.36$ Å) compared to MAPbI3 (with $Z_{\rm Pb}^{*}=4.59$, $Z_{\rm I}^{*}=4.44$, $\epsilon^{\infty}=5.87$, and $L=6.31$ Å) and CsPbI3 (with $Z_{\rm Pb}^{*}=4.49$, $Z_{\rm I}^{*}=4.17$, $\epsilon^{\infty}=5.75$, and $L=6.25$ Å). Nonetheless, our density functional perturbation theory calculations show that FAPbI3 exhibits the smallest LO-TO splitting among the three compounds. This contradiction indicates that the simplified expression does not capture the full microscopic picture Gonze and Lee (1997), as it neglects the phonon mode character, dipole directionality, and mode mixing. We ascribe the reduced LO-TO splitting in FAPbI3 to enhanced mode mixing between polar optical phonons and internal molecular vibrations [Fig. 14(c)], as well as directional coupling effects that diminish the effective dipole strength of the LO mode.
V.4 Overdamped phonon dynamics
@@
V.5 Thermal induced renormalization
We have seen that thermal induced corrections are strongly modified for most cubic compounds in the presence of local disorder. As shown in Fig. 10(a), electron-phonon contribution is strongly suppressed for all compounds, with the largest reductions found for Cs-based compounds which exhibit the largest degree of local disorder. The only case where electron-phonon coupling contributes negatively to the temperature coefficient of the band gap leading to a band gap closing is MAPbCl3. We attribute this to the fact that MAPbCl3 exhibits one of the highest degrees of local disorder among the hybrid halide perovskites, reaching levels comparable to those of Cs-based compounds [Fig. 2(b) and Fig. 2(c) of Ref. [Zacharias et al., 2026]]. The opposing effects of electron-phonon coupling and thermal expansion in polymorphous cubic MAPbCl3 lead to the nearly constant variation of the band gap with temperature, consistent with experiments.
@@ V.5 Thermal induced renormalization
Our results for the total temperature coefficients of the band gap using ZG polymorphous structures compare well with the experimental values, as shown quantitatively in Fig. 10(c) and Figs. 9(a)-(d), (i)-(l), and (q)-(t). Importantly, we have seen that even using a single ZG polymorphous configuration obtained from the MSS structure is a good approximation for capturing the temperature variation of the band gap. We also find that thermal-induced corrections lead to a linear variation of the effective masses with temperature [Figs. 9(e)-(h), (m)-(p), and (u)-(x)]. In APbI3 and APbBr3 compounds, the hole effective masses remain larger than the electron effective masses, whereas in ASnI3 and APbCl3 compounds, electron effective masses remain larger than hole effective masses. We note that our finite-temperature calculations: (i) rely on a constant temperature phonon approximation, (ii) neglect non-adiabatic effects in electron-phonon renormalization of the band structure Poncé et al. (2015); Miglio et al. (2020); Engel et al. (2022), (iii) neglect many-body quasiparticle corrections to the electronic structure Antonius et al. (2014); Marini et al. (2015); Li et al. (2024), and (iv) omit phonon renormalization due to thermal expansion. While including these effects may improve quantitative agreement with experiment, we do not expect them to alter the qualitative trends, as our current results already show good overall agreement.
@@
Figure 13: (a,b) Convergence test of the average DFT-PBEsol band gap of cubic FAPbI3 (a) and MAPbBr3 (b) as a function of the number of polymorphous configurations. (c) Convergence test of the average DFT-PBEsol band gap of cubic MAPbBr3 at 300 K [Eq. (11)] as a function of the number of ZG polymorphous configurations. Average values for 10 configurations are converged within 5 meV. (d) Convergence test of the average band gap temperature coefficient due to electron-phonon coupling as a function of the number of ZG polymorphous configurations used to compute the band gap of MAPbBr3 at each temperature. The temperature coefficient is calculated by fitting data points for the 5 temperatures shown in Fig. 7(c). The shaded regions represent the standard error of the mean on either side of the lines.
@@ VI.1 Electronic structure calculations
[…]
@@
PBE0 hybrid functional calculations Perdew et al. (1996); Adamo and Barone (1999) for the reference and polymorphous cubic structures were performed using the VASP code Kresse and Furthmüller (1996) within the projector augmented-wave formalism Kresse and Joubert (1999). A plane-wave energy cut-off of 400 eV was used, and the Brillouin zone was sampled with a $\times2$$\times$2 uniform k-grid.
[…]
@@ VI.3 Diffuse scattering
One-phonon diffuse scattering maps of MAPbBr3, shown in Fig. 6, were computed using disca.x of EPW Zacharias et al. (2021a, b); Lee et al. (2023). The Debye-Waller factors at each scattering wavevector $\bf Q$} were evaluated for a $\times30$$\times$30 $\bf q$}-grid using phonon polarization vectors and frequencies obtained by Fourier interpolation of the effective IFCs computed at 300 K with ASDM for $\times2$$\times$2 supercells. A $\times30$$\times$1 uniform $\bf Q$}-grid per Brillouin zone together with the computed Debye-Waller factors were used to calculate the one-phonon diffuse scattering intensity in reciprocal lattice planes perpendicular to the Cartesian $z$-axis. Parameters from Ref. [Peng et al., 2004] were used to model the atomic scattering amplitudes as a sum of Gaussians.
@@ VI.4 Effective masses
Effective masses were computed at the DFT level including SOC and by evaluating the diagonal components of the effective mass tensor via finite differences Giustino (2014), followed by an isotropic average. That is we used the second-order finite difference formula based on the energies at the R point and nearby $\bf k$}-points ($|\Delta{\bf k}|$ = 0.01 $2\pi$$\alpha$) along each Cartesian direction. The effective mass enhancement factors due to polymorphism were determined as $\lambda=m^{*}_{\rm poly}/m^{*}_{\rm ref}-1$ and due to thermal effects as $\lambda(T)=[m^{*}+\Delta m^{*}(T)]/m^{*}-1$. The reduced effective masses were evaluated as $\mu=m^{*}_{\rm h}m^{*}_{\rm e}/(m^{*}_{\rm h}+m^{*}_{\rm e})$. The slopes describing the temperature variation of the effective masses are denoted by $\lambda^{T}$.
@@ VI.5 Electron-phonon calculations
To account for electron-phonon coupling effects at various temperatures, we employed special displacements (or ZG displacements) Zacharias and Giustino (2020) generated via Eq. (6) for $\times4$$\times$4 supercells. We considered a constant temperature approximation for the phonons, i.e. anharmonic phonons computed at a specific temperature for each material (specified in Fig. 8), and the resulting vibrational modes were thermally populated at various temperatures to generate special displacements. We applied the special displacements on the nuclei of both the reference and polymorphous structures leading to the ZG reference and ZG polymorphous structures. We note that the initial $\times4$$\times$4 reference and polymorphous structures (containing 320 atoms for inorganic and 768 atoms for hybrid halide perovskites) were constructed by replicating the corresponding structures obtained in $\times2$$\times$2 supercells. SOC was included in all electron-phonon coupling calculations. To reduce computational cost, the plane wave energy cutoff was set to 60 Ry. Convergence tests performed on a subset of materials (MAPbBr3, FAPbBr3, FAPbI3, and MASnI3) using $\times2$$\times$2 supercells showed that the band gap values differ by less than 1 meV compared to calculations with a 120 Ry cutoff. Furthermore, the Brillouin zone of the $\times4$$\times$4 supercell was sampled using a $\times1$$\times$1 $\bf k$}-grid. In this case, we found for the same subset of materials that the band gap varied by less than 10 meV and the thermal induced renormalization by less than 1 meV. Figure 13(c) further shows that the band gap of MAPbBr3 at 300 K is already well converged with 6 ZG polymorphous configurations.
To determine the contribution of electron-phonon coupling to the band gap renormalization and effective masses ($dE_{\rm g}/dT|_{\rm EPC}$ and $dm^{*}/dT|_{\rm EPC}$), we applied linear fits to the data points of each material computed at various temperatures without considering the effect of thermal expansion. Each data point was obtained as a Boltzmann-weighted average [Eq. (11)] over the band gap of 10 ZG reference or 10 ZG polymorphous structures. Figure 13(d) shows the convergence of $dE_{\rm g}/dT|_{\rm EPC}$ computed for MAPbBr3 as a function of the number of ZG polymorphous configurations used to compute the band gap for each temperature.
[…]
@@ VI.6 Thermal expansion contribution
To evaluate the contribution of thermal expansion to the temperature dependence of the band gap, $dE_{\rm g}/dT|_{\rm TE}$, and effective masses, $dm^{*}/dT|_{\rm TE}$, we employed the following strategy for each material: (i) The lattice constants of the initial 10 reference and 10 polymorphous $\times2$$\times$2 supercells were varied by $\Delta L$ = $\pm$0.04 Å. (ii) For each modified lattice, geometry optimizations were performed while keeping the lattice constants fixed, yielding new reference and polymorphous configurations. (iii) Band gaps and effective masses were computed for each individual structure, and averaged over the set of configurations corresponding to the same lattice constant. (iv) The derivatives $dE_{\rm g}/dL$ and $dm^{*}/dL$ were estimated by fitting a linear function to the computed averaged values at $\Delta L, LL$, and $L+\Delta L$, equivalent to a central finite difference approximation. (v) Experimental values of the linear thermal expansion coefficient, $\alpha=1/L\,dL/dT$, of each material were taken from the literature Mashiyama et al. (2003); Rakita et al. (2015); Whitfield et al. (2016); Dang et al. (2016); Schueller et al. (2017); Kontos et al. (2018); Marronnier et al. (2018); Handa et al. (2020); He et al. (2021). Due to the absence of experimental data for FAPbCl3, we used $\alpha=0.6\times 10^{-4}$ K-1 close to the value reported for MAPbCl3 Handa et al. (2020). (vi) Finally, the thermal expansion contributions for each material were evaluated using the relations:
@@ VI.7 Phonon spectral functions
Phonon spectral functions of the polymorphous structures shown in Fig. 11 were obtained by phonon unfolding using the approach described in Ref. [Zacharias et al., 2023a] and the ZG.x code of EPW Lee et al. (2023). We used the IFCs calculated for the polymorphous structures without enforcing any translational or rotational symmetries of the unit cell or the reference structures. Based on the computed eigenvectors, which reflect spatial correlations introduced by local disorder, we determined the spectral weights of each configuration using 394 equally-spaced $\bf q$}-points along the the X-R-M-$\Gamma$-R path of the Brillouin zone. Convergence of the spectral weights was achieved by using a $\times12$$\times$12 $\bf g$}-grid of reciprocal lattice vectors.
@@ VII Conclusions and outlook
Based on the ASDM, we developed a unified methodology for electronic structure, anharmonicity, and electron-phonon coupling calculations in locally disordered inorganic and hybrid halide perovskites. We carried out a comparative high-throughput study of 12 cubic compounds and analyzed the modifications due to local disorder to their electronic structures with and without thermal effects. In addition to generating polymorphous structures, we also investigated reference configurations obtained by relaxing only the molecular sublattice, in order to minimize artifacts arising from net dipole moments, particularly relevant for MA-based compounds, and to enable systematic comparisons with the fully polymorphous structures. Our calculations reveal strong correlations between band gap widening and effective mass enhancement with structural distortions, such as changes in bond angles and lengths, induced by positional polymorphism. We further demonstrate that the extent of positional polymorphism is strongly influenced by the size of the A-site cation, a trend that correlates with the tolerance factor Zacharias et al. (2026), where lower values typically indicate increased structural flexibility and a greater propensity for local distortions. We also show the critical role of local disorder in explaining the experimentally observed band gap variations across different phases.
@@
We evaluated the temperature dependence of band gaps and effective masses using both ZG reference structures and ZG polymorphous structures for all cubic compounds, taking into account configurational entropy effects. For the band gaps, we achieved excellent agreement with experimental values, highlighting the importance of incorporating corrections due to positional polymorphism, exchange-correlation treatments beyond standard DFT, anharmonic electron-phonon coupling, and thermal expansion in order to accurately reproduce experimental trends. We further quantified the contributions of electron-phonon coupling and thermal expansion to the temperature-induced band gap renormalization, showing that electron-phonon coupling dominates in Cs-based and MA-based compounds, whereas thermal expansion plays the leading role in FA-based systems. This behavior is attributed to the larger volumetric thermal expansion coefficients of FA-based compounds. In the case of locally disordered FASnI3, we found that the stereochemically active Sn lone pair enhances the thermal expansion contribution, which primarily accounts for the largest temperature coefficient of the band gap among the materials studied. Finally, we showed that even a single ZG polymorphous configuration, generated from the MSS structure and ASDM phonons, yields results comparable to Boltzmann-weighted averages over 10 ZG polymorphous configurations.
@@
To our knowledge, this is the first report of anharmonic quasiparticle phonon dispersions in hybrid halide perovskites, validated against experimental measurements. The anharmonic, temperature-dependent interatomic force constants (IFCs), along with the corresponding reference and polymorphous structures, which are available in the NOMAD repository Zacharias (2025), provide a foundation for state-of-the-art perturbative electron-phonon calculations Baroni et al. (2001); Giustino (2017); Sangalli et al. (2019); Gonze et al. (2020); Engel et al. (2020); Zhou et al. (2021); Yang et al. (2022); Lee et al. (2023); Li et al. (2024) in cubic hybrid halide perovskites. These calculations can help elucidate key phenomena such as charge carrier mobilities Poncé et al. (2019); Ranalli et al. (2024), polaron formation Lafuente-Bartolome et al. (2024), as well as non-equilibrium dynamics Caruso (2021). Furthermore, our finding that a single ZG polymorphous configuration within a $4\times 4\times 4$ supercell provides an excellent approximation for nonperturbative electron-phonon calculations holds promise for extending this approach to large-scale, high-throughput studies across other hybrid halide perovskite systems.
@@
A database of monomorphous, reference, and polymorphous structures of halide perovskites, together with input and output files that led to the results of this study, is available in the NOMAD repository Zacharias (2025).